(b) (i) A random variable 'X' has the density function:


f(x) = K. — in- < x <


1
Expert's answer
2021-05-24T19:12:06-0400

(b) Probability density function of the random variable "X" is given by:


"f(x)=Kx(2-x)" , "\ 0<x<2"



We know that for "p.d.f" ,


"\int _{-\infty}^{\infty} f(x)dx=1"


"\Rightarrow \int _ {0}^{2} Kx(2-x)dx=1"


"\Rightarrow K\cdot\int _ {0}^{2}(2x-x^2)dx=1"


"\Rightarrow K\cdot |2.\frac{x^2}{2}-\frac {x^3}{3}|_{0}^{2}=1"


"\Rightarrow K\cdot[4-\frac{8}{3}]=1"


"\Rightarrow K=\dfrac{3}{4}\\"

Hence the value of K is "\boxed{K=\dfrac{3}{4}}"